# Towards Neural Scaling Laws for Time Series Foundation Models

### v1

@@ 3.2 Scaling Laws across Model Architectures
The above results suggest that the power-law captures the scaling behavior of the encoder-only Transformer in both ID and OOD scenarios. Similarly, we analyze the scaling properties of the decoder-only Transformer, along with two other state-of-the-art TSFMs, Chronos and Moirai, to assess the impact of model architectures on scaling behavior.
Encoder-only vs. Decoder-only Transformer. Fig. 5 presents a comparison of the scaling patterns between encoder-only and decoder-only Transformers. In terms of parameter scaling, the power law exponent of the encoder-only Transformer is consistently higher than that of the decoder-only Transformer across ID and OOD settings, indicating its superior scalability on parameters. Moreover, for a given number of parameters, the encoder-only Transformer generally achieves a lower log-likelihood loss, signifying better forecasting performance. Regarding compute scaling, both models demonstrate highly similar compute scalability on OOD data and slightly differ on ID data. When it comes to data, the scaling law fitting lines for encoder-only and decoder-only Transformers exhibit nearly constant offsets and similar slopes on data size scaling. This suggests that despite differences in performance, the scaling behavior of encoder-only and decoder-only Transformers is highly similar on pre-training dataset size.
Encoder-only Transformer vs. Moirai. Moirai is a TSFM based on an encoder-only Transformer architecture. It introduces any-variate attention to capture relationships between multiple variables. Additionally, Moirai incorporates a multi-scale patch embedding to handle different frequency patterns, and a more diverse mixture distribution to model real-world probability distributions. Fig. 6 shows a comparison between the scaling behavior of encoder-only Transformer and Moirai. On ID data, Moirai demonstrates better performance. However, for OOD data, as the number of model parameters increases, Moirai is gradually surpassed by the encoder-only Transformer. Comparing the power-law lines of the two models, we see that Moirai shows a smaller slope, indicating relatively weaker scalability. Collectively, Moirai shows significant improvements on ID time series forecast than our baseline, but some designs may limit its scalability on OOD data.
Decoder-only Transformer vs. Chronos. Chronos-T5 is an encoder-decoder Transformer-based TSFM that, like decoder-only Transformers, follows an auto-regressive prediction approach but uses a separate encoder to extract contextual information. It adopts point-wise prediction and transform numerical regression into discrete probability prediction. Fig. 7 compares the scaling patterns of decoder-only Transformers and Chronos-T5. When evaluated with log-likelihood, Chronos-T5 exhibits power laws with very small exponents, indicating limited scalability. We attribute this to the discrete probability prediction, as log-likelihood loss on a discrete distribution is not distance-sensitive, meaning the loss remains high unless the predicted value exactly matches the label. Compared to log-likelihood, symmetric mean absolute percentage error (SMAPE, definition is shown in Appendix C.3) is a more appropriate metric for evaluating the two models. Chronos-T5 shows a slight advantage in performance and scalability on ID data. However, on OOD data, the decoder-only Transformer performs much better. Despite some performance differences, both models exhibit similar scaling patterns on OOD data. Overall, the design improvements in Chronos-T5 enhance ID time series forecasting, but they do not effectively improve OOD generalization.
@@
Cross-Architecture Scaling Effects Summary. Our analysis of the scaling properties across architectures reveals several key findings: (1) Model architecture and design play a crucial role in determining scalability. (2) Encoder-only models demonstrate superior scalability and performance compared to decoder-only models, although both architectures exhibit similar scaling trends. (3) While Chronos and Moirai improve in-distribution ID forecasting, these gains do not extend effectively to OOD scenarios.
@@ 4 Design Principles for Time Series Foundation Models
Building on our findings regarding the scaling laws in TSFMs, we elaborate design principles to guide the development of effective and scalable models. These principles are framed around three key dimensions: training data, model parameters and architecture, as well as computational budget.
Training Data. Our experiments show that increasing the size of the training dataset leads to a greater performance improvement on OOD data compared to ID data. Enlarging the pre-training dataset is crucial for achieving better generalization. However, maintaining diversity within the dataset is equally important while increasing the data volume. Additionally, we observed that while there is a performance bias between encoder-only and decoder-only Transformers, they exhibit a consistent scaling pattern as the data size increases. This indicates that the benefits of data expansion are not influenced by model architecture, meaning that collecting more data can proceed in parallel with model improvements.
Model Parameters and Architecture. Our study highlights that model size is the most critical factor for improving TSFM performance. Among the three factors we examined, increasing the model size yielded the greatest benefit for ID forecast. In terms of architecture, an encoder-only Transformer architecture generally performs better and is more scalable than a decoder-only Transformer. Additionally, although the advanced models Chronos and Moirai showed some improvements in ID time series forecast, they failed to enhance OOD prediction capabilities. We speculate this is due to inductive biases introduced by the model modifications that do not generalize well. Compared to our baseline models, the scalability of Moirai and Chronos decreased, suggesting that the introduced designs limit the scalability of TSFMs. Model architecture significantly impacts scalability, and a good design should promote improvements in performance, generalization, and scalability.
Computational Budget. The power law we established indicates that there is an unbreakable lower bound for loss and MAPE under a given computational budget. This means that, with other factors remaining constant, as the model size increases, more compute resources must be invested to achieve better performance. However, different training objectives or model architectures can significantly affect this bound. For example, using masked recovery as a training objective typically shows a lower bound compared to next-token prediction. Similar to the scaling behavior of model size and dataset size, the increase in compute scale has a greater impact on improving OOD performance than on ID performance. This suggests that achieving robust generalization across distributions requires larger models, more data, and greater compute resources.
@@ 5 Related Works
Figure 8: Case studies on the “emergent abilities” of scaling time series model: We present three examples of zero-shot (unseen) out-of-distribution (OOD) time series prediction. In the US births, NN5 daily, and Australian electricity datasets, we observed that model behavior deviates from expected power law patterns, instead exhibiting characteristics more akin to emergent phenomena.
@@ 6.1 Additional Studies
Emergent Behaviors. Fig. 8 shows some examples of zero-shot OOD time series forecasting, where the model’s performance remains low until the model size reaches a critical threshold, after which performance improves significantly. This scaling behavior deviates from the previously observed power law and is more akin to emergent phenomena (Wei et al., 2022). This suggests that achieving good forecasting performance for certain time series may require a model with sufficient parameters to capture the underlying dynamic systems. For example, if a system of multistep differential equations requires $l$ sequential computation steps, the model may requires a depth of at least $O(l)$ layers to handle the complexity.
Sample Efficiency. Appendix D.2, Fig 9 shows the evaluation results during training. Large models are more sample-efficient than small models, reaching the better performance with the same optimization steps and using fewer time points.
Scaling Pattern Depends on Data Distributions. Appendix D.3, Fig. 10 shows the scaling behavior of TSFMs on Monash dataset, suggesting that OOD scaling behaviors depend on the relationship between the unseen data distribution and the training data distribution.
Scaling Pattern Depends on Performance Metrics. Appendix C.3, Fig. 11; 12;13 show the scaling behavior of continuous ranked probability score (CRPS), symmetric mean absolute percentage error(sMAPE), and mean absolute error (MAE). Comparing the results with log-likelihood and MAPE, we see obvious difference in the scaling patterns.
@@ 6.2 Conclusion
We have observed consistent scaling of encoder-only Transformer log-likelihood loss with parameter count, training computation and dataset size on both ID and OOD test data, as encapsulated in a power law. The experimental results show that as the number of parameters, computational resources, and training data increase, both models’ ID and OOD performance will continue to improve. Furthermore, our established laws suggest that for ID forecasting performance, larger model may be more important than more data. While, larger training dataset size plays a more crucial role in improving OOD forecasting performance.
We have also study the scaling properties of two common types of TSFMs, encoder-only Transformer and decoder-only Transformer. The two architectures show different scaling patterns in model size, computational resources, and training data. Generally, encoder-only Transformers show better performance and scalability. Moreover, we further include two advanced TSFMs, Chronos and Moirai, as specific cases to study the impact of model design on scaling behavior. We find that both Chronos and Moirai improve ID forecasting performance, but these gains can not translate effectively to OOD scenarios.
@@
Appendix A Appendix: Datasets and Setup Information
@@
Domain Balancing. In the Lotsa dataset, the climate domain accounts for nearly 90% of the data, while the remaining seven domains collectively represent only 10%. To create a more balanced pre-training dataset, we selected the traffic domain, which contains 4.8 billion time points, as the reference. We then scaled the data from the other domains to a similar magnitude.
Following the aforementioned processing steps, we compiled a high-quality time series dataset containing a total of 14.46 billion time points across 7 domains. The data volume is sufficient to support training models with parameters ranging from $10^{3}$ to $10^{8}$ parameters. Furthermore, each domain includes at least 100 million time points, ensuring that even when the dataset is divided into subsets of various sizes, domain diversity is preserved. Below, we introduce the datasets for each domain and outline the key properties of the datasets after processing, including domain type, sampling frequency, number of time series, total number of observations, and data source.
@@
Energy. The energy data primarily come from the BuildingsBench dataset, which provides time series data on residential and commercial building energy consumption. After applying quality filtering based on SNR, we selected samples from the Buildings-900K, BDG-2, and Sceaux datasets. Additionally, high-quality data from the ProEnFo library, including the GEF, ELF, PDB, Spanish, and Covid19_Energy datasets, were incorporated.
@@
Healthcare. We include a diverse sets of healthacre data from the ucr time series archive.
@@
Sales. Here we use the Favorita Sales dataset.
Table 2: Pretraining datasets and key properties.
@@
CMIP6 Climate 6H 8192 2,870,476,800 Nguyen et al. (2023)
ERA5 Climate H 8192 1,860,698,112 Nguyen et al. (2023)
Buildings900K Energy H 264980 2,321,489,780 Emami et al. (2023)
@@
BDG-2 Panther Energy H 136 893,840 Emami et al. (2023)
@@
Solar Power Energy 4S 26 5,248 Godahewa et al. (2021)
@@
Traffic Weekly Transport W 821 78,816 Godahewa et al. (2021)
Q-Traffic Transport 15T 45,148 257,200,384 Jiang et al. (2023)
PEMS04 Transport 5T 307 14,638,784 Jiang et al. (2023)
PEMS07 Transport 5T 883 23,789,760 Jiang et al. (2023)
PEMS08 Transport 5T 170 8,684,480 Jiang et al. (2023)
PEMS Bay Transport 5T 325 15,975,920 Jiang et al. (2023)
Loop Seattle Transport 5T 1,809 33,700,832 Jiang et al. (2023)
LargeST Transport 5T 8521 4,452,510,528 Liu et al. (2023)
Azure VM Traces 2017 CloudOTS 5T 159,472 880,648,165 Woo et al. (2023)
Borg Cluster Data 2011 CloudOTS 5T 143,386 176,650,715 Woo et al. (2023)
Alibaba Cluster Trace 2018 CloudOTS 5T 58409 191,740,824 Woo et al. (2023)
@@
Kaggle Web Traffic Weekly Web W 145,063 15,206,232 Godahewa et al. (2021)
@@ Appendix B Time Series Models
Patch Embedding. We split the input $(\mathbf{x}_{1},\dots,\mathbf{x}_{L})$ into $M$ univariate time series $\mathbf{x}^{(i)}\in\mathbb{R}^{1\times L}$, independently forecasting future time series for each variate. Each univariate time series $\mathbf{x}^{(i)}$ is first divided into non-overlapped patches. Specifically, given the patch length as $P$, the patching process will generate the a sequence of patches $\mathbf{x}^{(i)}_{p}\in\mathbb{R}^{P\times N}$ where $N=\left[\frac{L}{P}\right]$ is the number of patches. Then the patches are mapped to the latent space of $d_{m}$ via a learnable linear projection $\mathbf{W}_{p}\in\mathbb{R}^{d_{m}\times P}$. In our baseline models, the patch size $P$ is set to 32.
@@
is the rotary matrix with pre-defined paramters $\Theta=\{\theta_{i}=10000^{-2(i-1)/d},i\in[1,2,\dots,d/2]\}$, $\mathbf{W}_{q,k}$ is the learned query or key projection weights, and $\mathbf{x}_{m}\in\mathbb{R}^{d}$ is the embedding of the $m$ token. Next, applying RoPE to the dot product of query and key, we can obtain:
@@
where $\phi(\cdot),\varphi(\cdot)$ are usually non-negative functions, e.g. elu($\cdot$)+1, or exp($\cdot$). Overall, unlike the traditional approach of adding positional information before the query and key projection, RoPE utilizes a rotary matrix to transform query and key embeddings, effectively leveraging the geometric properties of vectors
@@
with parameters $\nu>0,\mu\in\mathbb{R},\tau>0$ represent the degrees of freedom (df), location, and scale coefficient, respectively, and $\Gamma$ is the gamma function. We modeled the predicted distribution using a mixture of four Student-t distributions. Our baseline model employs independent linear layers to predict the distribution parameters: degrees of freedom, location, scale, and mixture weights. To ensure positive values for the scale and df, we apply a soft-plus transformation. The mixture weights are constrained within the probability simplex using a soft-max function. Additionally, to avoid undefined variance for low df values, we impose a lower bound of 2 on the df parameter
Encoder-only Transformer. The encoder-only Transformer has proven effective for time series tasks (Nie et al., 2023). In input time seires, the future horizon is replaced by learnable mask tokens combined with position information. After passing through bidirectional attention blocks, the future representations are derived from these mask tokens and mapped to the parameters of a mixture distribution. Through sampling from the mixture distribution, we can get predictions for future patches. The pseudo code of encoder-only Transformer-based TSFM is as follows:
{minted}
[frame=single,framesep=5pt]python class EncoderOnlyTransformer(nn.Module): patch_tokenizer = nn.Linear(N, d_m) encoder = TransformerEncoder() mask_token_embedding = nn.Embedding(1, d_m) mixture_distribution_head = MixtureDistributionHead()
def forecast(self, x, T): B, L, N = x.shape # Patch Embedding: (B, L, N)-¿(B, L, d_m) x = patch_tokenizer(x)
# Add mask tokens: (B, L, d_m)-¿(B, L+T, d_m) mask_tokens = self.mask_token_embedding.weight.expand(B, T, -1) x = torch.cat((x, mask_tokens), dim=1)
# Pass through bidirectional attention blocks encoded_output = self.encoder(x)
# Truncate the future tokens: (B, L+T, d_m)-¿(B, T, d_m) future_tokens = encoded_output[:,-self.num_mask_tokens:]
# Predict the parameters of mixture distribution future_dist = self.mixture_distribution_head(future_tokens)
return future_dist
Decoder-only Transformer. The key difference in the decoder-only Transformer is its unidirectional attention. In the input sequence, no masking is needed. The future horizon is predicted based on the preceding token. In other words, given a sequence of input patches, the model is optimized to predict the next patch as a function of all past patches. Similar to LLMs this can be done in parallel over the entire context window, and automatically enables the model to predict the future after having seen varying number of input patches. The pseudo code of decoder-only Transformer-based TSFM is as follows:
{minted}
[frame=single,framesep=5pt]python class DecoderOnlyTransformer(nn.Module): patch_tokenizer = nn.Linear(N, d_m) decoder = nn.TransformerDecoder() mixture_distribution_head = MixtureDistributionHead()
def forecast(self, x, T): # Patch Embedding: (B, L, N)-¿(B, L, d_m) x = self.patch_tokenizer(x)
# Generate predictions for future time steps step-by-step preds = []
for t in range(T): # If it’s the first loop, use the original input tgt = torch.cat([x,*preds],dim=1) if preds else x
# Pass through unidirectional attention blocks decoded_output = self.decoder(tgt)
# Take the last token next_token = decoded_output[:, -1:, :]
# Record the prediction of the current time step predictions.append(next_token)
# Concatenate all predictions: (B, T, d_m) future_tokens = torch.cat(predictions, dim=1)
# Predict the parameters of mixture distribution future_dist = self.mixture_distribution_head(future_tokens)
return future_dist

### v2

@@ 3.2 Scaling Laws across Model Architectures
The above results suggest that the power-law captures the scaling behavior of encoder-only Transformers in both ID and OOD scenarios. Similarly, we analyze the scaling properties of decoder-only Transformers, along with two other state-of-the-art TSFMs, Chronos and Moirai, to assess the impact of model architectures on scaling behavior. These models are trained on the dataset we built with the same training setup.
Encoder-only vs. Decoder-only Transformer. Figure 5 presents a comparison of the scaling patterns between encoder-only and decoder-only Transformers. In parameter scaling, the encoder-only Transformer shows a slight advantage, with a marginally higher power-law exponent on ID data and lower NLL values. For compute and data scaling, the two architectures demonstrate nearly identical scalability across ID and OOD settings, with some performance differences observed on ID data. Overall, the two architectures exhibit similar scalability, aside from a performance difference observed in ID data.
@@
Encoder-only Transformer vs. Moirai. Moirai is a TSFM based on an encoder-only Transformer architecture. It introduces ”any-variate” attention to capture relationships between multiple variables. Additionally, Moirai incorporates a multi-scale patch embedding to handle different frequency patterns, and a more diverse mixture distribution to model real-world probability distributions. Figure 6 show a comparison between the scaling behavior of encoder-only Transformer and Moirai. On ID data, Moirai demonstrates better performance. However, for OOD data, as the number of model parameters increases, Moirai is gradually surpassed by the encoder-only Transformer. Comparing the power-law lines of the two models, we see that Moirai shows a smaller slope, indicating relatively weaker scalability. Collectively, Moirai shows significant improvements on ID time series forecast than our baseline, but some designs may limit its scalability on OOD data.
@@
Decoder-only Transformer vs. Chronos. Chronos-T5 is an encoder-decoder Transformer-based TSFM that, like decoder-only Transformers, follows an auto-regressive prediction approach but uses a separate encoder to extract contextual information. It adopts point-wise prediction and transform numerical regression into discrete probability prediction. Figure 7 compares the scaling patterns of decoder-only Transformers and Chronos-T5. When evaluated with NLL, Chronos-T5 exhibits power laws with very small exponents, indicating limited scalability. We attribute this to the discrete probability prediction, as NLL on a discrete distribution is not distance-sensitive, meaning the loss remains high unless the predicted value exactly matches the label. Compared to NLL, symmetric mean absolute percentage error (SMAPE, definition is shown in Appendix C.3) is a more appropriate metric for evaluating the two models. Chronos-T5 shows a slight advantage in performance and scalability on ID data. However, on OOD data, the decoder-only Transformer performs better. Despite some performance differences, both models exhibit similar scaling patterns on OOD data. Overall, the design improvements in Chronos-T5 enhance ID time series forecasting, but they do not effectively improve OOD generalization.
Scaling Effects Summary. Our analysis of the scaling properties across architectures reveals several key findings: (1) Model architecture and design play a crucial role in determining scalability. (2) Encoder-only models and decoder-only models show a similar scalability on OOD data; while encoder-only models have a slightly higher parameter scalability and a certain performance advantage on ID data. (3) While Chronos and Moirai improve ID forecasting, these gains do not extend effectively to OOD scenarios.
@@ 4 Design Principles for Time Series Foundation Models
Building on our findings regarding the scaling laws in TSFMs, we elaborate design principles to guide the development of scalable models. These principles are framed around three key dimensions: training data, model parameters and architecture, as well as computational resources.
Training Data. Our experiments show that increasing the size of the training dataset leads to a greater performance improvement on OOD data compared to ID data. Enlarging the pre-training dataset is crucial for achieving better generalization. Based on Equation 4, we can expect that doubling the size of the pre-training dataset will result in an OOD MAPE reduction to approximately 90% of its previous value, and an ID MAPE reduction to approximately 97%. But maintaining diversity within the dataset is equally important while increasing the data volume.
Model Parameters and Architecture. Our study highlights that model size is the most critical factor for improving TSFM performance. Compared to collecting more data, increasing model size yields the greater benefit for both ID and OOD forecast. However, as model size scales up, data bottlenecks inevitably emerge. From the ratio of exponents in Equations 2 and 4, we infer that scaling up model size should be accompanied by a sublinear increase in dataset size. Specifically, this relationship can be approximated as $D\propto N^{\frac{\alpha_{N}}{\alpha_{D}}}\sim N^{0.8}$, meaning that doubling the model size requires roughly 1.7 times more data. In terms of model architecture, we found some modifications facilitate ID performance but can not generalize well to OOD scenarios. Model design can impact scalability, and a good design should promote performance, generalization, and scalability.
Compute. The power law we established indicates that there is an unbreakable lower bound for NLL loss and MAPE under a given computational budget. This means that, with other factors remaining constant, as the model size increases, more compute resources must be invested to achieve better performance. Our findings suggest that compute resources should prioritize scaling up model size rather than extending training time for fixed models as larger models are more sample-efficient than smaller models. Moreover, different training objectives or model architectures can significantly affect this performance bound, exploring a compute-efficient training paradigm will be promising.
Figure 8: Case studies on the “Emergent Abilities” of scaling time series models. We present three examples of zero-shot (unseen) out-of-distribution (OOD) time series prediction. In the us births, nn5 daily, and australian electricity datasets, we observed that model behavior deviates from expected power law patterns, instead exhibiting characteristics more akin to emergent phenomena.
@@ 6.1 Additional Studies
Emergent Behaviors. Figure 8 shows some examples of zero-shot OOD time series prediction, where the model’s performance remains low until the model size reaches a critical threshold, after which performance improves significantly. This scaling behavior deviates from the previously observed power law and is more akin to emergent phenomena (Wei et al., 2022).
Sample Efficiency. Appendix D.2, Figure 14 shows the evaluation results during training. Large models are more sample-efficient than small models, reaching better performance with the same optimization steps and using fewer time points.
Scaling Pattern Depends on Data Distributions. Appendix D.3, Figure 15 - 19 show the scaling behavior of TSFMs on the Monash subset, suggesting that OOD scaling behaviors varies depending on the relationship between the unseen data distribution and the pre-training data distribution.
Scaling Pattern Depends on Performance Metrics. Appendix D.4, Figure 15 - 19 provides a comprehensive analysis of scaling behavior across five performance metrics: NLL, MAPE, SMAPE, MASE, and CRPS. All metrics exhibit a consistent decreasing trend, approximately following a power law.
@@ 6.2 Conclusion
We have observed consistent scaling of encoder-only Transformer NLL with parameter count, training computation and dataset size on both ID and OOD test data, as encapsulated in a power law. The experimental results show that as the number of parameters, computational resources, and training data increase, both models’ ID and OOD performance will continue to improve. Furthermore, our established laws suggest that for ID and OOD forecasting performance, larger model may be more important than more data.
We have also study the scaling properties of two common types of TSFMs, encoder-only Transformer and decoder-only Transformer. The two architectures show similar scalability in OOD data across model size, computational resources, and training data. While encoder-only models have a slightly better parameter scalability and performance in ID data based our evaluation setting. Moreover, we further include two advanced TSFMs, Chronos and Moirai, as specific cases to study the impact of module design on scaling behavior. We find that both Chronos and Moirai improve ID forecasting performance, but these gains can not translate effectively to OOD scenarios. Overall, model architecture and module design is important influence factor for scalability.
@@
Acknowledgments and Disclosure of Funding
This research was partly supported by Australian Research Council (ARC) under grants FT210100097 and DP240101547 and the CSIRO – National Science Foundation (US) AI Research Collaboration Program.
Ethics Statement
Our research is dedicated exclusively to addressing scientific challenges and does not involve human participants, animals, or materials that pose environmental concerns. We anticipate no ethical risks or conflicts of interest.
Reproducibility Statement
We provide the implementation details in Appendices B and C, including model configurations and training & evaluation details. The source code and related source of this work are available at https://github.com/Qingrenn/TSFM-ScalingLaws for reproducibility.
Appendix A Pre-training Datasets
@@
Domain Balancing. In the LOTSA dataset, the climate domain accounts for nearly 90% of the data, while the remaining seven domains collectively represent only 10%. To create a more balanced pre-training dataset, we selected the transport domain, which contains 4.8 billion time points, as the reference. We then scaled the data from the other domains to a similar magnitude.
Following the aforementioned processing steps, we compiled a high-quality time series dataset containing a total of 16.8 billion time points across 7 domains. The data volume is sufficient to support training models with parameters ranging from $10^{3}$ to $10^{8}$ parameters. Furthermore, most domains include at least 100 million time points, ensuring that even when the dataset is divided into subsets of various sizes, domain diversity is preserved. Below, we introduce the datasets for each domain and outline the key properties of the datasets after processing, including domain type, sampling frequency, number of time series, total number of observations, and data source.
@@
Energy. The energy data primarily come from the BuildingsBench dataset, which provides time series data on residential and commercial building energy consumption. After applying quality filtering based on SNR, we selected samples from the Buildings-900K, BDG-2, and Sceaux datasets. Additionally, high-quality data from the ProEnFo library, including the GEF, ELF, PDB, Spanish, and Covid19__\__Energy datasets, were incorporated.
@@
Healthcare. We include a diverse sets of healthcare data from the ucr time series archive.
@@
Sales. Here we use the Favorite Sales dataset.
Table 2: Pre-training datasets and key properties.
@@
CMIP6 Climate 6H 196,608 2,870,476,800 Nguyen et al. (2023)
ERA5 Climate H 212,992 1,860,698,112 Nguyen et al. (2023)
Buildings900K Energy H 538,577 4,718,473,097 Emami et al. (2023)
@@
BDG-2 Panther Energy H 105 919,800 Emami et al. (2023)
@@
Solar Power Energy 4S 1 7,397,222 Godahewa et al. (2021)
@@
Traffic Weekly Transport W 862 82,752 Godahewa et al. (2021)
Q-Traffic Transport 15T 45,148 264,386,688 Jiang et al. (2023)
PEMS04 Transport 5T 921 15,649,632 Jiang et al. (2023)
PEMS07 Transport 5T 883 24,921,792 Jiang et al. (2023)
PEMS08 Transport 5T 510 9,106,560 Jiang et al. (2023)
PEMS Bay Transport 5T 325 16,941,600 Jiang et al. (2023)
Loop Seattle Transport 5T 1,809 33,953,760 Jiang et al. (2023)
LargeST Transport 5T 42,333 4,452,510,528 Liu et al. (2023)
Azure VM Traces 2017 CloudOTS 5T 159,472 885,522,908 Woo et al. (2023)
Borg Cluster Data 2011 CloudOTS 5T 286,772 1,075,105,708 Woo et al. (2023)
Alibaba Cluster Trace 2018 CloudOTS 5T 116,818 190,385,060 Woo et al. (2023)
@@
Kaggle Web Traffic Weekly Web W 145,063 16,537,182 Godahewa et al. (2021)
@@ Appendix B Time Series Models
Patch Embedding. We split the input $(\mathbf{x}_{1},\dots,\mathbf{x}_{L})$ into $M$ uni-variate time series $\mathbf{x}^{(i)}\in\mathbb{R}^{1\times L}$, independently forecasting future time series for each variate. Each uni-variate time series $\mathbf{x}^{(i)}$ is first divided into non-overlapped patches. Specifically, given the patch length as $P$, the patching process will generate the a sequence of patches $\mathbf{x}^{(i)}_{p}\in\mathbb{R}^{P\times N}$ where $N=\left[\frac{L}{P}\right]$ is the number of patches. Then the patches are mapped to the latent space of $d_{m}$ via a learnable linear projection $\mathbf{W}_{p}\in\mathbb{R}^{d_{m}\times P}$. In our baseline models, the patch size $P$ is set to 32.
@@
is the rotary matrix with pre-defined parameters $\Theta=\{\theta_{i}=10000^{-2(i-1)/d},i\in[1,2,\dots,d/2]\}$, $\mathbf{W}_{q,k}$ is the learned query or key projection weights, and $\mathbf{x}_{m}\in\mathbb{R}^{d}$ is the embedding of the $m$ token. Next, applying RoPE to the dot product of query and key, we can obtain:
@@
where $\phi(\cdot),\varphi(\cdot)$ are usually non-negative functions, e.g. elu($\cdot$)+1, or exp($\cdot$). Overall, unlike the traditional approach of adding positional information before the query and key projection, RoPE utilizes a rotary matrix to transform query and key embeddings, effectively leveraging the geometric properties of vectors.
Figure 9: Gaussian mixture (GM) distribution vs. Student-t mixture (STM) distribution. Left: GM generally under-performs compared to STM in both ID and OOD test data, particularly as model size increases. Right: GM exhibits unstable loss convergence during training, while STM achieves smoother convergence. These findings underscore the prevalence of long-tail distributions in real-world time series data and highlights the crucial role of Student-t mixtures in effectively modeling these distributions.
@@
with parameters $\nu>0,\mu\in\mathbb{R},\tau>0$ represent the degrees of freedom (df), location, and scale coefficient, respectively, and $\Gamma$ is the gamma function. We modeled the predicted distribution using a mixture of four Student-t distributions. Our baseline model employs independent linear layers to predict the distribution parameters: degrees of freedom, location, scale, and mixture weights. To ensure positive values for the scale and df, we apply a soft-plus transformation. The mixture weights are constrained within the probability simplex using a soft-max function. To avoid undefined variance for low df values, we impose a lower bound of 2 on the df parameter.
Additionally, we conducted an experiment to compare the performance of Gaussian mixture distribution and Student-t mixture distribution when used as prior distributions to approximate real-world time series. Figure 9 shows that Student-t mixtures generally outperform Gaussian mixtures, particularly as model size increases. Furthermore, we observe unstable convergence when using Gaussian mixture distribution as the prediction head. These findings highlight the prevalence of long-tail distributions in real-world time series data and emphasize the critical role of Student-t mixtures in effectively modeling such distributions.
Encoder-only Transformer. The encoder-only Transformer has proven effective for time series tasks (Nie et al., 2023). In input time series, the future horizon is replaced by learnable mask tokens combined with position information. After passing through bidirectional attention blocks, the future representations are derived from these mask tokens and mapped to the parameters of a mixture distribution. Through sampling from the mixture distribution, we can get predictions for future patches.
Decoder-only Transformer. The key difference in the decoder-only Transformer is its unidirectional attention. In the input sequence, no masking is needed. The future horizon is predicted based on the preceding token. In other words, given a sequence of input patches, the model is optimized to predict the next patch as a function of all past patches. Similar to LLMs this can be done in parallel over the entire context window, and automatically enables the model to predict the future after having seen varying number of input patches.